# Logarithmic Finite-Size Scaling of the Four-Dimensional Ising Model

### v1

@@ Abstract
The Ising model is one of the most fundamental models in statistical physics and plays a crucial role in facilitating the comprehensive analysis of phase transitions and critical phenomena duminil2022100. For a lattice $G=(V,E)$ with the vertex set $V$ and edge set $E$, the Hamiltonian of the Ising model reads
@@
where $s_{i}\in\{-1,+1\}$ denotes the spin on the $i$-th vertex and the summation runs over all edges on the lattice, $h$ refers to the external magnetic field and $J>0$ is the coupling strength. The partition function is then given by $\mathcal{Z}=\sum_{s}e^{-\beta\mathcal{H}(s)}$. Let $K:=\beta J$ be the reduced coupling strength. Hereinafter, we set $J=1$ and focus on the zero-field case with $h=0$ in this paper. If the spin is extended to be an $n$-component vector with unit length, then Eq. (1) is the Hamiltonian of the O($n$) model ONmodel, where the case $n=1$ is the Ising model.
In most cases, the Ising model cannot be exactly solved, making the investigation of its critical behaviors heavily reliant on numerical methods on finite systems, such as Monte Carlo (MC) simulations. To analyze the results, the finite-size scaling (FSS) method is employed as a powerful method, which describes the asymptotic approach of finite systems to the thermodynamic limit near a continuous phase transition point $K_{c}$, to effectively estimate critical points and exponents SuzukiFSS; FSSPriv; brezin1985finite; nishimori2011elements. The main assumption of FSS is that the correlation length is effectively truncated by the linear system size $L$, such that the singular part of the free energy density function for a $d$-dimensional system can be written as
@@
Above the upper critical dimension, $d>d_{c}=4$, the scaling behaviors of the Ising model are characterized by the RG exponents given by mean field theory PhysRevLett.47.1, or, specifically, by the Gaussian fixed point (GFP) as $(y_{t},y_{h})=(2,1+d/2)$. However, if one considers the system with periodic boundary condition (PBC), such standard FSS breakdowns brezin1985finite; BDofFSS; FSSforhighDIsingtori. For example, it is observed that $\chi\sim L^{d/2}$ in the PBC case, which is different from $\chi\sim L^{2}$ as the standard FSS in Eq. (4) predicts. Recent numerical and theoretical results suggest that the scaling form of the free energy function, as described by Eq. (2), is conjectured to have an extended form as 5DFK2020
@@
where $(y_{t}^{*},y_{h}^{*})=(d/2,3d/4)$ are obtained from the exact calculation of the complete-graph (CG) Ising model  [^1] A complete graph with V vertices is a graph on which each vertex is connected to all others, which can be seen as the application of the Landau mean field theory to finite systems luijten1997interaction. The $\tilde{f}_{0}$ term corresponds to the GFP asymptotics, accounts for spatial fluctuations and governs the FSS of distance-dependent observables. The $\tilde{f}_{1}$ term represents the CG asymptotics and governs the leading FSS of various macroscopic observables, such as $\chi$ and $c_{\textsc{e}}$.
At $d_{c}=4$, the two sets of mean-field RG exponents coincide with each other, i.e., $(y_{t},y_{h})=(y^{*}_{t},y^{*}_{h})$. Field theory predicts that at the upper critical dimension, multiplicative logarithmic corrections appear. For the O$(n)$ model, in the thermodynamic limit, when approaching the criticality $(t\to 0)$, susceptibility and specific heat are predicted to diverge as larkin1969phase; RGanalysis1973
@@
In terms of the FSS, the singular part of the finite-size free-energy density involving the multiplicative logarithmic corrections for the O$(n)$ model is proposed in Ref. kenna2004finite as
@@
with $\hat{y}_{t}=(4-n)/(2n+16),\hat{y}_{h}=1/4$. The $n=1$ case, which is the Ising model, is studied in Ref. aktekin2001finite where $(\hat{y}_{t},\hat{y}_{h})=(1/6,1/4)$. Recently, as inspired from the high-dimensional scaling form (Eq. (6)), the FSS of the free energy density at $d_{c}=4$ is conjectured in Ref. lv2021finite as
@@
Compared with Eq. (9), the key feature in Eq. (10)-(11) is the simultaneous existence of the GFP and the CG (after modified with multiplicative logarithmic corrections) asymptotics in the FSS formula of the free energy density, and the logarithmic corrections only apply to the CG term, not to the GFP term. Accordingly, the leading FSS of various macroscopic quantities suffer from logarithmic corrections, while quantities purely controlled by the GFP, such as the short-distance decay of $g(\mathbf{r},L)$, are free from logarithmic corrections.
Specifically, for the 4D Ising model at the critical point, from Eqs. (10) and (11), one can obtain the FSS behaviors for the following quantities:
@@
However, numerically verifying the logarithmic corrections is a challenging task. Earlier in 1987, a logarithmic scaling behavior of the specific heat was roughly observed in the 4D Ising model from a Monte Carlo (MC) study in Ref. ESanchez-Velasco_1987. However, since the simulated system size was not large enough, the result was inconclusive. Recently, a systematic numerical study in Ref. lv2021finite has been done to verify the multiplicative logarithmic corrections described by Eqs. (10) and (11) in the 4D Ising, XY and Heisenberg models, respectively corresponding to $n=1,2,3$ case of the O$(n)$ model. The FSS behaviors of $\chi$, $\chi_{\mathbf{k}}$ and $g(\mathbf{r},L)$ are observed to be consistent with the predictions from Eqs. (10) and (11), providing strong evidence to the existence of the magnetic correction exponent $\hat{y}_{h}$. However, the expected logarithmic divergence of the specific heat has not been clearly observed in Ref. lv2021finite. Later on, in Ref. 4DSAW, the authors numerically studied the logarithmic FSS for the 4D self-avoiding walk(SAW) model, which corresponds to the $n\rightarrow 0$ case of the O$(n)$ model, and both the exponents $\hat{y}_{t}=\hat{y}_{h}=1/4$ were clearly observed. So, it turns out that observing the thermal correction exponent $\hat{y}_{t}$ for the Ising model is much more challenging. In fact, in Refs. lundow2023revising; lundow2009Ising, the authors suggest that the specific heat of the Ising model is bounded, i.e., $\hat{y}_{t}=0$. Meanwhile, a large-size simulation up to $L=1024$ in Ref. HOTRGfor4DIsing using higher-order tensor RG method also fails to detect the multiplicative logarithmic correction in the specific heat.
Besides the conventional spin representation, the Ising model can also be effectively described and analyzed using two geometric representations: the Fortuin-Kasteleyn (FK) bond representation and the loop representation fortuin1972random; van1941lange, which will be called the FK Ising model and the loop Ising model. These representations are derived from certain expansions of the partition function, offering alternative perspectives for studying the properties of the Ising model. Specifically, the FK Ising model is the $q=2$ case of the general $q$-state random-cluster (RC) model. Given a graph $\mathcal{G}$, each edge of the graph is either occupied by a bond or not. Then, the partition function of the $q$-state RC model is
@@
where $w=\tanh{K}$ is the weight of each occupied bond and the Kronecker delta function $\delta_{F\in{\rm even}(G)}$ serves as an indicator function that ensures that for any subgraph $F$ giving a nonzero contribution to the partition function is an even graph. Such representation is also called the random-current model, random even graph or the flow representation of the Ising model HD2016RC; GrimmiteRandomEvenGraph.
@@
Recently, many numerical studies in Refs. HighDFK1; HighDFK2; 5DFK2020; FKCG2021; LoopCG2023; xiao2023finitesize have been conducted on the two geometric representations of the Ising model above $d_{c}=4$ and on the CG (which can be regarded as the $d\to\infty$ limit). Compared to the spin Ising model, the FK and loop Ising models exhibit much richer geometric properties, such as two length scales, two configuration sectors and two scaling windows. In particular, numerical results strongly suggest there simultaneously exist two upper critical dimensions ($d_{c}=4,d_{p}=6$) in the FK Ising model; one can refer to Fig. 17 in Ref. HighDFK1 and Tabel I in Ref xiao2023finitesize for the summary of main results. Generally speaking, the scaling behavior of the largest FK cluster and large loop clusters (size $\gg O(L^{2})$) are controlled by the CG Ising asymptotics, while other FK clusters and medium loop clusters (size $\leq O(L^{2})$) are described by the GFP asymptotics. Moreover, a multiplicative logarithmic correction has been discovered in the scaling of the second-largest FK cluster as $C_{2}\sim L^{y_{h}}(\ln{L})^{-1}$ for all $d>4$, which has not been found by any theoretical investigation.
@@
Inspired by these insightful results, we systematically explore the logarithmic corrections to various geometric quantities of the 4D FK and loop Ising models, in particular to check whether the thermal correction exponent $\hat{y}_{t}$ or the expected divergence of the specific heat can be clearly observed in geometric representations. We use Wolff and Swendsen-Wang algorithms to simulate the FK Ising model, and the lifted worm algorithm to simulate the loop Ising model. We study the scaling of the susceptibility $\chi$ in the FK and loop Ising models, and both suggest that $\chi\sim L^{2}(\ln{L})^{1/2}$, consistent to the theoretical prediction and previous numerical results in Ref. lv2021finite. Moreover, we also observe that the largest FK cluster $C_{1}$, which is a magnetic quantity, scales as $L^{3}(\ln{L})^{\hat{y}_{h}}$ with $\hat{y}_{h}=1/4$. The sizes of loop clusters in the loop Ising model are energy-like quantities. Since it is conjectured in Eq. (10) that the logarithmic corrections only apply to the modified CG term and the large loop clusters are believed to follow the CG-loop-Ising asymptotics, we hope the large loop clusters suffer much less additive finite-size corrections from the GFP term compared with energy quantities in the spin Ising model. Indeed, our results suggest that the size of the largest loop cluster scales as $F_{1}\sim L^{y_{t}}(\ln{L})^{\hat{y}_{t}}$ with $y_{t}=2$ and $\hat{y}_{t}=1/6$. In addition, we study the variance of the number of loop bonds $c_{\textsc{b}}$, which is shown to exhibit the same leading FSS as the spin specific heat, and our data strongly suggest $c_{\textsc{b}}\sim(\ln{L})^{2\hat{y}_{t}}$. Thus, the long-standing subtle question on the scaling of the specific heat of the spin Ising model is clarified in the loop Ising model.
@@
Additionally, we study the size of the second largest cluster $C_{2}$ in the FK Ising model and our data suggest that $C_{2}\sim L^{3}(\ln{L})^{-1/4}$, implying a new magnetic logarithmic correction exponent $\hat{y}_{h2}=-1/4$, which has no theoretical prediction. Therefore, even in terms of logarithmic corrections, the Ising model under geometric representations also exhibit richer phenomena than the spin Ising model.
@@
Meanwhile, we employ the lifted worm algorithmWormAlg; LiftedWorm to generate the loop configuration and sample the following observables in the loop representation
@@
    The average returning time $T_{w}=\langle\mathcal{T}_{w}\rangle$, which is equivalent to the spin susceptibility $\chi$ in the spin representation WormAlg;
@@ III.1 Magnetic scaling behaviors
In this section, we first discuss the scaling behaviors of some magnetic quantities in geometric representations. It can be shown that both the second moment of sizes of all FK-clusters $S_{2}$ and the returning time $T_{w}$ correspond to the spin susceptibility $\chi$ PhysRevD.38.2009; grimmett2006random; PhysRevLett.87.160601.
@@
Next, we examine the effect of logarithmic corrections on the size of clusters in the FK Ising model. It is expected that the fractal dimension of the FK clusters is equal to the magnetic exponent $y_{h}$, as numerically observed above 4D in Refs. HighDFK1; HighDFK2. Therefore, one can expect that at 4D, the power-law scaling behavior of the FK clusters is dominated by $L^{y_{h}}$ with $y_{h}=3$. Likewise, we introduce the rescaled cluster sizes $\tilde{C}_{n}:=C_{n}/L^{3}(n=1,2)$ for the largest and second-largest clusters $C_{1}$ and $C_{2}$. In Fig. 1(b1), we plot in log-log scale $\tilde{C}_{1}$ versus $\ln{L/L_{0}}$ with $L_{0}$ fixed at $0.55$. The data collapse nicely onto a straight line with slope $1/4$, indicating that $C_{1}\sim L^{3}[\ln(L/L_{0})]^{1/4}$. Similar to Fig. 1(a2), we plot $\tilde{C}_{1}^{4}$ against $\ln{L}$ in Fig. 1(b2), to remove the uncertainty from the constant $L_{0}$. Again, the nice linear relationship implies that $C_{1}\sim L^{3}({a_{1}\ln L+a_{2}})^{1/4}$. So at 4D, our data suggest that $C_{1}\sim L^{y_{h}}({\ln L})^{\hat{y}_{h}}$ with $\hat{y}_{h}=1/4$.
Then we investigate the scaling behavior of the second-largest cluster $C_{2}$. In the high-d FK Ising model ($d>4$), a multiplicative logarithmic correction has been observed for the second-largest cluster, which scales as $C_{2}\sim L^{1+d/2}({\ln L})^{-1}$ HighDFK1. This indicates at 4D, the second-largest cluster may exhibit different logarithmic corrections to the largest cluster. Figure 2(1) plots $\tilde{C}_{2}$ versus $\ln{L/L_{0}}$ in the log-log scale, with $L_{0}$ fixed at $1.73$. As one can see, the correction term decreases as the system size increases, which is totally different from $C_{1}$ shown in Fig. 1(b1). The data collapse onto the dashed line with slope $-1/4$ in the figure suggests that $C_{2}\sim L^{3}(\ln{L})^{-1/4}$. Furthermore, we plot $1/\tilde{C}_{2}^{4}$ versus $\ln{L}$ in Fig. 2(2), and the data points collapse well onto a straight line. This confirms the aforementioned scaling behaviour for $C_{2}$, i.e., the logarithmic correction in $C_{2}$ is characterized by a new exponent $\hat{y}_{h2}=-1/4$. This finding is particularly noteworthy as the logarithmic correction in $C_{2}$ that has not been previously identified in either theoretical or numerical studies at $d_{c}=4$. We note that this result is also different from the high-d cases ($d>d_{c}$), where $\hat{y}_{h2}=-1$.

### v2

@@ Abstract
The Ising model is one of the most fundamental models in statistical physics and plays a crucial role in facilitating the comprehensive analysis of phase transitions and critical phenomena Duminil-Copin (2022). For a lattice $G=(V,E)$ with the vertex set $V$ and edge set $E$, the Hamiltonian of the Ising model reads
@@
where $s_{i}\in\{-1,+1\}$ denotes the spin on the $i$-th vertex and the summation runs over all edges on the lattice, $h$ refers to the external magnetic field and $J>0$ is the coupling strength. The partition function is then given by $\mathcal{Z}=\sum_{s}e^{-\beta\mathcal{H}(s)}$. Let $K:=\beta J$ be the reduced coupling strength. Hereinafter, we set $J=1$ and focus on the zero-field case with $h=0$ in this paper. If the spin is extended to be an $n$-component vector with unit length, then Eq. (1) is the Hamiltonian of the O($n$) model Stanley (1968), where the case $n=1$ is the Ising model.
In most cases, the Ising model cannot be exactly solved, making the investigation of its critical behaviors heavily reliant on numerical methods on finite systems, such as Monte Carlo (MC) simulations. To analyze the results, the finite-size scaling (FSS) method is employed as a powerful method, which describes the asymptotic approach of finite systems to the thermodynamic limit near a continuous phase transition point $K_{c}$, to effectively estimate critical points and exponents Suzuki (1977); Privman (1990); Brézin and Zinn-Justin (1985); Nishimori and Ortiz (2011). The main assumption of FSS is that the correlation length is effectively truncated by the linear system size $L$, such that the singular part of the free energy density function for a $d$-dimensional system can be written as
@@
Above the upper critical dimension, $d>d_{c}=4$, the scaling behaviors of the Ising model are characterized by the RG exponents given by mean field theory Aizenman (1981), or, specifically, by the Gaussian fixed point (GFP) as $(y_{t},y_{h})=(2,1+d/2)$. However, if one considers the system with periodic boundary condition (PBC), such standard FSS breakdowns Brézin and Zinn-Justin (1985); Grimm et al. (2017); Zhou et al. (2018). For example, it is observed that $\chi\sim L^{d/2}$ in the PBC case, which is different from $\chi\sim L^{2}$ as the standard FSS in Eq. (4) predicts. Recent numerical and theoretical results suggest that the scaling form of the free energy function, as described by Eq. (2), is conjectured to have an extended form as Fang et al. (2020)
@@
where $(y_{t}^{*},y_{h}^{*})=(d/2,3d/4)$ are obtained from the exact calculation of the complete-graph (CG) Ising model  [^1] A complete graph with V vertices is a graph on which each vertex is connected to all others, which can be seen as the application of the Landau mean field theory to finite systems Luijten (1997). The $\tilde{f}_{0}$ term corresponds to the GFP asymptotics, accounts for spatial fluctuations and governs the FSS of distance-dependent observables. The $\tilde{f}_{1}$ term represents the CG asymptotics and governs the leading FSS of various macroscopic observables, such as $\chi$ and $c_{\textsc{e}}$.
At $d_{c}=4$, the two sets of mean-field RG exponents coincide with each other, i.e., $(y_{t},y_{h})=(y^{*}_{t},y^{*}_{h})$. Field theory predicts that at the upper critical dimension, multiplicative logarithmic corrections appear. For the O$(n)$ model, in the thermodynamic limit, when approaching the criticality $(t\to 0)$, susceptibility and specific heat are predicted to diverge as Larkin and Khmel’Nitskiĭ (1969); Wegner and Riedel (1973)
@@
In terms of the FSS, the singular part of the finite-size free-energy density involving the multiplicative logarithmic corrections for the O$(n)$ model is proposed as
@@
with $\hat{y}_{t}=(4-n)/(2n+16),\hat{y}_{h}=1/4$ Kenna (2004); Kenna et al. (2006a, b). The $n=1$ case, which is the Ising model, is studied in Ref. Aktekin (2001) where $(\hat{y}_{t},\hat{y}_{h})=(1/6,1/4)$. Recently, as inspired from the high-dimensional scaling form (Eq. (6)), the FSS of the free energy density at $d_{c}=4$ is conjectured in Ref. Lv et al. (2021) as
@@
Compared with Eq. (9), the key feature in Eq. (10)-(11) is the simultaneous existence of the GFP and the CG (after modified with multiplicative logarithmic corrections) asymptotics in the FSS formula of the free energy density, and the logarithmic corrections only apply to the CG term, not to the GFP term. Accordingly, the leading FSS of various macroscopic quantities suffer from logarithmic corrections, while quantities purely controlled by the GFP, such as the short-distance decay of $g(\mathbf{r},L)$, are free from logarithmic corrections. Specifically, for the 4D Ising model at the critical point, from Eqs. (10) and (11), one can obtain the FSS behaviors for the following quantities:
@@
However, numerically verifying the logarithmic corrections is a challenging task. Earlier in 1987, a logarithmic scaling behavior of the specific heat was roughly observed in the 4D Ising model from a Monte Carlo (MC) study in Ref. Sanchez-Velasco (1987). However, since the simulated system size was not large enough, the result was inconclusive. Recently, a systematic numerical study in Ref. Lv et al. (2021) has been done to verify the multiplicative logarithmic corrections described by Eqs. (10) and (11) in the 4D Ising, XY and Heisenberg models, respectively corresponding to $n=1,2,3$ case of the O$(n)$ model. The FSS behaviors of $\chi$, $\chi_{\mathbf{k}}$ and $g(\mathbf{r},L)$ are observed to be consistent with the predictions from Eqs. (10) and (11), providing strong evidence to the existence of the magnetic correction exponent $\hat{y}_{h}$. However, the expected logarithmic divergence of the specific heat has not been clearly observed in Ref. Lv et al. (2021). Later on, in Ref. Fang et al. (2021a), the authors numerically studied the logarithmic FSS for the 4D self-avoiding walk(SAW) model, which corresponds to the $n\rightarrow 0$ case of the O$(n)$ model, and both the exponents $\hat{y}_{t}=\hat{y}_{h}=1/4$ were clearly observed. So, it turns out that observing the thermal correction exponent $\hat{y}_{t}$ for the Ising model is much more challenging. In fact, in Refs. Lundow and Markström (2023, 2009), the authors suggest that the specific heat of the Ising model is bounded, i.e., $\hat{y}_{t}=0$. Meanwhile, a large-size simulation up to $L=1024$ in Ref. Akiyama et al. (2019) using higher-order tensor RG method also fails to detect the multiplicative logarithmic correction in the specific heat.
Besides the conventional spin representation, the Ising model can also be effectively described and analyzed using two geometric representations: the Fortuin-Kasteleyn (FK) bond representation and the loop representation Fortuin and Kasteleyn (1972); van der Waerden (1941), which will be called the FK Ising model and the loop Ising model. These representations are derived from certain expansions of the partition function, offering alternative perspectives for studying the properties of the Ising model. Specifically, the FK Ising model is the $q=2$ case of the general $q$-state random-cluster (RC) model. Given a graph $\mathcal{G}$, each edge of the graph is either occupied by a bond or not. Then, the partition function of the $q$-state RC model is
@@
where $w=\tanh{K}$ is the weight of each occupied bond and the Kronecker delta function $\delta_{F\in{\rm even}(G)}$ serves as an indicator function that ensures that for any subgraph $F$ giving a nonzero contribution to the partition function is an even graph. Such representation is also called the random-current model, random even graph or the flow representation of the Ising model Duminil-Copin (2018); Grimmett and Janson (2007).
@@
Recently, many numerical studies in Refs. Fang et al. (2023, 2022, 2020, 2021b); Li et al. (2023); Xiao et al. (2024) have been conducted on the two geometric representations of the Ising model above $d_{c}=4$ and on the CG (which can be regarded as the $d\to\infty$ limit). Compared to the spin Ising model, the FK and loop Ising models exhibit much richer geometric properties, such as two length scales, two configuration sectors and two scaling windows. In particular, numerical results strongly suggest there simultaneously exist two upper critical dimensions ($d_{c}=4,d_{p}=6$) in the FK Ising model; one can refer to Fig. 17 in Ref. Fang et al. (2023) and Tabel I in Ref Xiao et al. (2024) for the summary of main results. Generally speaking, the scaling behavior of the largest FK cluster and large loop clusters (size $\gg O(L^{2})$) are controlled by the CG Ising asymptotics, while other FK clusters and medium loop clusters (size $\leq O(L^{2})$) are described by the GFP asymptotics. Moreover, a multiplicative logarithmic correction has been discovered in the scaling of the second-largest FK cluster as $C_{2}\sim L^{y_{h}}(\ln{L})^{-1}$ for all $d>4$, which has not been found by any theoretical investigation.
@@
Inspired by these insightful results, we systematically explore the logarithmic corrections to various geometric quantities of the 4D FK and loop Ising models with PBC, in particular to check whether the thermal correction exponent $\hat{y}_{t}$ or the expected divergence of the specific heat can be clearly observed in geometric representations. We use Wolff and Swendsen-Wang algorithms to simulate the FK Ising model, and the lifted worm algorithm to simulate the loop Ising model. We study the scaling of the susceptibility $\chi$ in the FK and loop Ising models, and both suggest that $\chi\sim L^{2}(\ln{L})^{1/2}$, consistent to the theoretical prediction and previous numerical results in Ref. Lv et al. (2021). Moreover, we also observe that the largest FK cluster $C_{1}$, which is a magnetic quantity, scales as $L^{3}(\ln{L})^{\hat{y}_{h}}$ with $\hat{y}_{h}=1/4$. The sizes of loop clusters in the loop Ising model are energy-like quantities. Since it is conjectured in Eq. (10) that the logarithmic corrections only apply to the modified CG term and the large loop clusters are believed to follow the CG-loop-Ising asymptotics, we hope the large loop clusters suffer much less additive finite-size corrections from the GFP term compared with energy quantities in the spin Ising model. Indeed, our results suggest that the size of the largest loop cluster scales as $F_{1}\sim L^{y_{t}}(\ln{L})^{\hat{y}_{t}}$ with $y_{t}=2$ and $\hat{y}_{t}=1/6$. In addition, we study the variance of the number of loop bonds $c_{\textsc{b}}$, which is shown to exhibit the same leading FSS as the spin specific heat, and our data strongly suggest $c_{\textsc{b}}\sim(\ln{L})^{2\hat{y}_{t}}$. Thus, the long-standing subtle question on the scaling of the specific heat of the spin Ising model is clarified in the loop Ising model.
@@
Additionally, we study the size of the second-largest cluster $C_{2}$ in the FK Ising model and our data suggest that $C_{2}\sim L^{3}(\ln{L})^{-1/4}$, implying a new magnetic logarithmic correction exponent $\hat{y}_{h2}=-1/4$, which has no theoretical prediction. Therefore, even in terms of logarithmic corrections, the Ising model under geometric representations also exhibit richer phenomena than the spin Ising model.
@@
Meanwhile, we employ the lifted worm algorithmProkof’ev and Svistunov (2001a); Elçi et al. (2018) to generate the loop configuration and sample the following observables in the loop representation
@@
    The average returning time $T_{w}=\langle\mathcal{T}_{w}\rangle$, which is equivalent to the spin susceptibility $\chi$ in the spin representation Prokof’ev and Svistunov (2001a);
@@
Here we employ a pseudorandom number generator based on the modulo-2 addition of two independent shift registers with lengths chosen as the Mersenne exponents 127 and 9689. This generator is well tested in Ref. Shchur and Blöte (1997), and no biased error has been found thus far.
@@ III.1 Magnetic scaling behaviors
In this section, we first discuss the scaling behaviors of some magnetic quantities in geometric representations. It can be shown that both the second moment of sizes of all FK-clusters $S_{2}$ and the returning time $T_{w}$ correspond to the spin susceptibility $\chi$ Edwards and Sokal (1988); Grimmett (2006); Prokof’ev and Svistunov (2001b).
@@
Next, we examine the effect of logarithmic corrections on the size of clusters in the FK Ising model. It is expected that the fractal dimension of the FK clusters is equal to the magnetic exponent $y_{h}$, as numerically observed above 4D in Refs. Fang et al. (2023, 2022). Therefore, one can expect that at 4D, the power-law scaling behavior of the FK clusters is dominated by $L^{y_{h}}$ with $y_{h}=3$. Likewise, we introduce the rescaled cluster sizes $\tilde{C}_{n}:=C_{n}/L^{3}(n=1,2)$ for the largest and second-largest clusters $C_{1}$ and $C_{2}$. In Fig. 1(b1), we plot in log-log scale $\tilde{C}_{1}$ versus $\ln{L/L_{0}}$ with $L_{0}$ fixed at $0.55$. The data collapse nicely onto a straight line with slope $1/4$, indicating that $C_{1}\sim L^{3}[\ln(L/L_{0})]^{1/4}$. Similar to Fig. 1(a2), we plot $\tilde{C}_{1}^{4}$ against $\ln{L}$ in Fig. 1(b2), to remove the uncertainty from the constant $L_{0}$. Again, the nice linear relationship implies that $C_{1}\sim L^{3}({a_{1}\ln L+a_{2}})^{1/4}$. So at 4D, our data suggest that $C_{1}\sim L^{y_{h}}({\ln L})^{\hat{y}_{h}}$ with $\hat{y}_{h}=1/4$.
Then we investigate the scaling behavior of the second-largest cluster $C_{2}$. In the high-d FK Ising model ($d>4$), a multiplicative logarithmic correction has been observed for the second-largest cluster, which scales as $C_{2}\sim L^{1+d/2}({\ln L})^{-1}$ Fang et al. (2023). This indicates at 4D, the second-largest cluster may exhibit different logarithmic corrections to the largest cluster. Figure 2(1) plots $\tilde{C}_{2}$ versus $\ln{L/L_{0}}$ in the log-log scale, with $L_{0}$ fixed at $1.73$. As one can see, the correction term decreases as the system size increases, which is totally different from $C_{1}$ shown in Fig. 1(b1). The data collapse onto the dashed line with slope $-1/4$ in the figure suggests that $C_{2}\sim L^{3}(\ln{L})^{-1/4}$. Furthermore, we plot $1/\tilde{C}_{2}^{4}$ versus $\ln{L}$ in Fig. 2(2), and the data points collapse well onto a straight line. This confirms the aforementioned scaling behaviour for $C_{2}$, i.e., the logarithmic correction in $C_{2}$ is characterized by a new exponent $\hat{y}_{h2}=-1/4$. This finding is particularly noteworthy as the logarithmic correction in $C_{2}$ that has not been previously identified in either theoretical or numerical studies at $d_{c}=4$. We note that this result is also different from the high-d cases ($d>d_{c}$), where $\hat{y}_{h2}=-1$.